# THE RESTRICTED PICARD FUNCTOR

### v2

@@ 1 Introduction
Representability results for Picard varieties are therefore of fundamental importance. A wonderful theorem, proved by Murre [Mur] and Oort [Oor] building on work of Grothendieck [FGA], states that the Picard functor $\operatorname{Pic}_{X/K}$ of a proper scheme over a field $K$ is represented by a scheme [Mur, Th. 2]. Beyond the proper case one cannot expect such representability, as one already sees in the case of the affine line. Indeed, the Picard group of the affine line over a field is trivial, so were $\operatorname{Pic}_{\mathbf{A}^{1}/K}$ represented by a scheme (necessarily locally of finite type), it would have to be infinitesimal. But for a reduced scheme $X$ that is not seminormal, one does not generally have $\operatorname{Pic}(\mathbf{A}^{1}_{X})=\operatorname{Pic}(X)$, so the desired representability cannot hold. A better, more conceptual obstacle to representability is that a simple cohomological calculation using the ring of dual numbers as a test scheme shows that the tangent space to $\operatorname{Pic}_{X/K}$ at the identity is given by ${\rm{H}}^{1}(X,\mathcal{O}_{X})$, which will not be finite in general beyond the proper setting.
One observes that the problem in the above examples arises from test schemes that are not “nice.” Thus Achet, following a suggestion of Raynaud, was led to introduce the restricted Picard functor – which is the Picard functor but with one’s test schemes restricted to be smooth over $K$ – and proved its representability for forms of affine space admitting a regular compactification [Ach, Th. 1.1].
@@
Unfortunately, even the restricted Picard functor cannot be represented by a scheme in general. To illustrate the problem, let $\overline{C}/K$ be a smooth projective geometrically connected curve of genus $g>0$, let $x,y\in\overline{C}(K)$ be distinct rational points, and let $C:=\overline{C}\backslash\{x,y\}$. Then $\operatorname{Pic}_{\overline{C}/K}$ is an extension of $\mathbf{Z}$ by an abelian variety, the Jacobian $J$ of $\overline{C}$. Because $\operatorname{Pic}(\overline{C}\times T)\twoheadrightarrow\operatorname{Pic}(C\times T)$ for every smooth (or even regular) $K$-scheme $T$, one can conclude that $\operatorname{Pic}^{+}_{C/K}$ is the quotient of $J$ by the point $j$ represented by the divisor $[x]-[y]$. If $j$ is not a torsion point, then this quotient is not a scheme. Thus we see that an inevitable obstacle to representability comes from quotients of group schemes by subgroups generated by points of infinite order. We may now state the main result of the present paper. For a commutative, locally finite type $K$-group scheme $G$, we denote by $G^{t}$ the preimage in $G$ of the torsion subgroup of the étale component group $G/G^{0}$.
@@
If $K$ is a field and $X$ is a regular $K$-scheme of finite type, then there is an isomorphism $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$, where $G$ is a smooth commutative $K$-group scheme with finitely generated component group and $E\subset G^{t}$ is an étale $K$-group scheme such $E(K_{s})$ is a free abelian group of finite rank. Furthermore, $G^{0}$ is unique up to unique isomorphism.
For results on the structure of the group $G^{0}$ appearing in Theorem 1.3, see §6. As we shall see, the most technically demanding case of Theorem 1.3 is when $K$ is imperfect. Indeed, for perfect $K$, one may entirely skip §§34.
Theorem 1.3 is nice, but what we would really like are results asserting representability by a scheme. Using the above theorem, however, it will not be hard to deduce the following.
@@
Let $K$ be a field and $X$ a regular $K$-scheme of finite type with dense smooth locus. Assume that any map from $X_{K_{s}}$ into an abelian variety $A/K_{s}$ factors through a finite subscheme of $A$. Then $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth commutative $K$-group scheme with finitely generated component group. Furthermore, $(\operatorname{Pic}^{+}_{X/K})^{0}$ is wound unipotent.
@@
Because abelian varieties do not admit nonconstant maps from rationally connected varieties, $X$ admits no nonconstant map to an abelian variety over $\overline{K}$, hence also not over $K_{s}$. Now apply Theorem 1.4. ∎
@@
Among the examples falling under the rubric of Theorem 1.6 are smooth connected unipotent groups. These may indeed have positive-dimensional restricted Picard functors when $K$ is imperfect. For examples, see for instance [Ros2, Props. 5.2, 5.4, 5.5, Ex. 5.8].
@@
The problem, of course, is that we do not know whether $X$ admits a regular compactification. Even when $K$ is algebraically closed, in positive characteristic the existence of a regular compactification remains a formidable open problem. Fortunately, however, there is a weaker form of resolution which is adequate for many purposes: de Jong’s alterations theorem [deJ1, Th. 4.1]. (Actually, we will require a somewhat more precise result from [deJ2].) The difference between an alteration and a resolution is that we relax the requirement that the resolution be a birational map and only require that it be generically finite. We may replace $X$ by a dense open subscheme and thereby assume that the alteration $X^{\prime}\rightarrow X$ is finite flat. We then use faithfully flat descent to describe a line bundle on $X$ in terms of descent datum for a line bundle on $X^{\prime}$. This is helpful because $X^{\prime}$ admits a regular compactification and therefore has “almost-representable” restricted Picard functor in the sense of Theorem 1.3.
The key difficulty in showing that the functor whose points consist of a line bundle plus descent datum on $X^{\prime}$ is (almost) representable is to prove that the descent data which arise in this manner are “uniformly bounded” in a suitable sense. To illustrate the difficulty, consider the situation in which $X=\mathbf{A}^{1}_{K}$ and $X^{\prime}=\mathbf{A}^{1}_{K^{\prime}}$ for some finite extension field $K^{\prime}/K$. (Of course, this is in a sense a silly example because $\operatorname{Pic}^{+}_{X/K}=0$ in this case, as one may see directly. But it illustrates the point most simply – without some of the technical window dressing – and so the author hopes that the reader will indulge him and pretend that he does not realize this already.) A line bundle on $X^{\prime}$ is then automatically trivial, and a descent datum for $X^{\prime}_{T}/X_{T}$ consists of nothing other than a unit on $(X^{\prime}\times_{X}X^{\prime})_{T}=(\mathbf{A}^{1}_{T})_{K^{\prime}\otimes_{K}K^{\prime}}$ (which is to say, an automorphism of the trivial line bundle) that satisfies the cocycle condition. The problem is that $K^{\prime}\otimes_{K}K^{\prime}$ may well be non-reduced. Indeed, in the situation of main interest for our result – when $K$ is imperfect – this will typically be the case. Therefore a unit on $(\mathbf{A}^{1}_{K^{\prime}\otimes_{K}K^{\prime}})_{T}$ may have arbitrarily high degree in the variable $y$ on $\mathbf{A}^{1}$. Without bounding this degree a priori, we would obtain not representability by a scheme, but by an ind-scheme, because the set of all polynomials is not represented by a scheme. Thus we require some sort of bound on the degree. This will follow from the cocycle condition, which says that our unit $u$ satisfies
@@
The structure of this paper is as follows. In §2 we descend the almost-representability of the restricted Picard functor through finite Galois covers. In §3 we discuss the notion of boundedness for a collection of descent data, and show how boundedness of the collection of all possible descent data on certain line bundles implies the almost-representability of the restricted Picard functor. In §4 we use the notion of boundedness to construct the tools required to descend almost-representability of the restricted Picard functor through purely inseparable extension fields of degree $p$. In §5 we prove the almost-representability of the restricted Picard functor for regular $K$-schemes of finite type. Finally, in §6, we study the structure of the restricted Picard functor.
@@ 2 Descent through Galois covers
Let us first prove the following claim. Given a normal $K$-scheme $X$ of finite type, a dense open subscheme $U\subset X$, and an element $a\in\Gamma(U,\mathcal{O}_{U})$ algebraic over $K$, $a$ extends uniquely to a section of $X$. Indeed, let $p\in X\backslash U$ be a codimension $1$ point. Then $a|U$ defines a rational map $g\colon X\dashrightarrow\mathbf{P}^{1}_{K}$ which necessarily extends over $p$. Because $a$ is algebraic over $K$, $g|U$ lands inside a finite closed subscheme $Z$ of $\mathbf{A}^{1}_{K}$, and $g^{-1}(Z)$ therefore defines a closed subscheme of a neighborhood $W$ of $p$ containing $W\cap U$, hence all of $W$. We have thus extended $a$ over a neighborhood of $p$, and the reducedness of $W$ ensures that this extension is unique. Thus we obtain a unique extension of $a$ over each codimension $1$ point of $X$, and because $X$ is normal, this then extends to all of $X$, as claimed.
Now to the assertion of the proposition. The uniqueness of ${\rm{Units}}_{X/K}$ follows from Yoneda’s Lemma, as does the fact that it is a commutative $K$-group scheme. To prove existence and the other assertions, we are free by Galois descent to replace $K$ by a finite Galois extension. First assume that $X$ is separated, and choose a normal compactification $\overline{X}$ of $X$. Let $A:=\Gamma(\overline{X},\mathcal{O}_{\overline{X}})$, a finite $K$-algebra, hence algebraic over $K$. Note that its image in $\Gamma(U,\mathcal{O}_{U})$ consists exactly of the elements algebraic over $K$, thanks to the claim of the first paragraph. We have an inclusion $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})\hookrightarrow f_{*}{\mathbf{G}}_{\rm{m}}$, and we wish to show among other things that this map induces an isomorphism onto ${\rm{Units}}_{X/K}^{0}$. For this we are free to replace $K$ by a finite Galois extension. Extending scalars, we may assume that all of the boundary components $E_{1},\dots,E_{n}$ of $\overline{X}\backslash X$ are geometrically irreducible. For any smooth $K$-scheme $T$, $X_{T}$ is still normal, and a uniformizer of $D_{i}$ is also a uniformizer of $(D_{i})_{T}$. We thus obtain a homomorphism
@@
given by taking valuations along each of the $D_{i}$, where $f_{\operatorname{\acute{e}t}}$ is the map on étale sites induced by $f$. Then $\ker(\phi)(T)=(\overline{f}_{*}{\mathbf{G}}_{\rm{m}})(T)$ for every smooth $T/K$, where $\overline{f}\colon\overline{X}\rightarrow{\rm{Spec}}(K)$ is the structural morphism. If $A:=\Gamma(\overline{X},\mathcal{O}_{\overline{X}})$, a finite $K$-algebra, then $\overline{f}_{*}{\mathbf{G}}_{\rm{m}}=\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$. Replacing $K$ by a finite Galois extension, we may assume that $\operatorname{im}(\phi)$ is constant and that every $K$-point in $\operatorname{im}(\phi)$ lifts to a point of $(f_{\operatorname{\acute{e}t}})_{*}{\mathbf{G}}_{\rm{m}}$. So there exist units $u_{1},\dots,u_{n}$ on $X$ whose images freely generate $\operatorname{im}(\phi)$. Then every unit on $X_{T}$ for smooth $T$ is in the free subgroup generated by the $u_{i}$ times a unit extending over $\overline{X}_{T}$, which is to say, an element of $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})(T)$. Thus we obtain a smooth $K$-group scheme ${\rm{Units}}_{X/K}$ which is an extension of a free finite rank constant group scheme by $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$, which computes the units of $X_{T}$ for smooth $T$, and which is a subfunctor of the functor of units (because $u_{i}$ and points of $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$ define units on $X_{T}$ for any $K$-scheme $T$). This completes the proof when $X$ is separated.
Now consider the general, possibly non-separated case. Let $U\subset X$ be a separated dense open subscheme (for instance, a dense affine open). Replacing $K$ by a finite Galois extension if necessary, we may assume that all of the boundary components of $X\backslash U$ are geometrically irreducible. Let $D_{1},\dots,D_{m}$ be these components. For any smooth scheme $T$, $X_{T}$ is still normal, and a uniformizer of $D_{i}$ is also a uniformizer of $(D_{i})_{T}$. We thus obtain a homomorphism
@@
given by taking valuations along each of the $D_{i}$, and $\ker(\phi)(T)=(f_{*}{\mathbf{G}}_{\rm{m}})(T)$ for every smooth $T/K$. Because the codomain of $\phi$ is étale, $\ker(\phi)^{0}={\rm{Units}}_{U/K}^{0}$ and $\ker(\phi)$ is smooth. Furthermore, the algebraic subrings of $\Gamma(X,\mathcal{O}_{X})$ and $\Gamma(U,\mathcal{O}_{U})$ agree by the first paragraph. Thus $\ker(\phi)^{0}=\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$, where $A\subset\Gamma(X,\mathcal{O}_{X})$ is a finite $K$-algebra, hence $\ker(\phi)^{0}$ defines a subfunctor of $f_{*}{\mathbf{G}}_{\rm{m}}$. A similar argument to the one in the first paragraph involving the group freely generated by the $u_{i}$ then shows that $\ker(\phi)$ is a subfunctor of $f_{*}{\mathbf{G}}_{\rm{m}}$, hence ${\rm{Units}}_{X/K}=\ker(\phi)$. That the component group is free of finite rank follows from the analogous assertion for $U$. ∎
@@
For a smooth commutative $K$-group scheme $M$ of finite type equipped with an action by the finite group $G$, $\mathscr{H}^{i}(G,M)$ is represented by a smooth commutative $K$-group scheme of finite type, which is affine if $M$ is so.
@@
The functor $\mathcal{Z}^{n}(G,M)$ of $n$-cocycles $G^{n}\rightarrow M$ is readily checked to be representable by the closed subscheme of $M^{G^{n}}$ corresponding to the cocycle condition. We also have coboundary maps $d_{n+1}\colon M^{G^{n}}\rightarrow\mathcal{Z}^{n}(G,M)$. Because $M^{G^{n}}$ is smooth, the quotient group sheaf $Q^{i}:=\mathcal{Z}^{i}(G,M)/{\rm{im}}(d_{i+1})$ on the smooth-étale site of $K$ agrees with the usual fppf quotient group scheme. Replacing $Q^{i}$ by its maximal smooth $K$-subgroup scheme ([CGP, Lem.C.4.1, Rem. C.4.2]) does not change the underlying functor on the smooth-étale site of $K$, so we see that $\mathscr{H}^{i}(G,M)$ is represented by a smooth commutative finite type $K$-group scheme, which is affine if $M$ is. ∎
@@
Let $M$ be a smooth commutative $K$-group scheme with finitely generated component group, and suppose that $M$ comes equipped with an action by a finite group $G$. Then ${\mathscr{H}}^{i}(G,M)$ is (represented by) a smooth $K$-group scheme with finitely generated component group. For $i>0$, ${\mathscr{H}}^{i}(G,M)$ is finite type, and it is affine if $M^{0}$ is so.
@@
Let $E:=M/M^{t}$ be the torsion-free quotient of the component group of $M$. By Galois descent, we may assume that $K$ is separably closed, so that $E$ is constant and the quotient map $M\rightarrow E$ admits a section, so that $M\simeq M^{t}\times E$. We are thus reduced to the étale and finite type cases, which are handled by Propositions 2.3 and 2.4. ∎
@@
We say that a sheaf $\mathscr{F}$ of abelian groups on the smooth-étale site of ${\rm{Spec}}(K)$ is almost-representable when it is of the form $G/E$ for a smooth commutative $K$-group scheme $G$ with finitely generated component group and a subsheaf $E\subset G^{t}$ represented by an étale $K$-group scheme $E$ with $E(K_{s})$ free of finite rank.
@@
We will need the following lemma in order to recognize almost-representable sheaves.
@@
If $G$ is a smooth commutative $K$-group scheme with finitely generated component group, and $E\subset G$ is a subsheaf on the smooth-étale site of ${\rm{Spec}}(K)$ represented by an étale $K$-group scheme with $E(K_{s})$ finitely generated, then $G/E$ is almost-representable.
@@
Let $C:=G/G^{t}$ denote the torsion-free quotient of the component group of $G$, and let $\phi\colon E\rightarrow C$ denote the composition $E\subset G\rightarrow C$. First assume that $\phi(E)=0$. Then $E\subset G^{t}$. Let $T\subset E$ denote the torsion subgroup. Then $\overline{G}:=G/T$ is a smooth $K$-group scheme with finitely-generated component group, and $G/E=\overline{G}/(E/T)$, so $G/E$ is almost-representable.
Now assume that $\phi(E)\subset C$ is nonzero. Then $E$ defines an element of $\operatorname{Ext}^{1}(\phi(E),\ker(\phi))$. By Lemma 2.8, this element is killed by some $N>0$. The multiplication by $N$ map induces an isomorphism $\phi(E)\simeq[N]\phi(E)$, so the extension

### v4

@@ 1 Introduction
Representability results for Picard functors are therefore of fundamental importance. A wonderful theorem, proved by Murre [Mur] and Oort [Oor] building on work of Grothendieck [FGA], states that the Picard functor $\operatorname{Pic}_{X/K}$ of a proper scheme over a field $K$ is represented by a scheme [Mur, Th. 2]. Beyond the proper case one cannot expect such representability, as one already sees in the case of the affine line. Indeed, the Picard group of the affine line over a field is trivial, so were $\operatorname{Pic}_{\mathbf{A}^{1}/K}$ represented by a scheme (necessarily locally of finite type), it would have to be infinitesimal. But for a reduced scheme $X$ that is not seminormal, one does not generally have $\operatorname{Pic}(\mathbf{A}^{1}_{X})=\operatorname{Pic}(X)$, so the desired representability cannot hold. A better, more conceptual obstacle to representability is that a simple cohomological calculation using the ring of dual numbers as a test scheme shows that the tangent space to $\operatorname{Pic}_{X/K}$ at the identity is given by ${\rm{H}}^{1}(X,\mathcal{O}_{X})$, which will not be finite-dimensional in general beyond the proper setting.
One observes that the problem in the above examples arises from test schemes that are not “nice.” Thus Achet, following a suggestion of Raynaud, was led to introduce the restricted Picard functor – which is the Picard functor but with one’s test schemes restricted to be smooth over $K$ – and proved its representability for forms of affine space admitting a regular compactification [Ach, Th. 1.1]. (By a form $X$ of affine $n$-space we mean an fppf form. By the Nullstellensatz, this is the same as saying that $X_{L}\simeq\mathbf{A}^{n}_{L}$ as $L$-schemes for some finite field extension $L/K$.)
@@
Unfortunately, even the restricted Picard functor cannot be represented by a scheme in general. To illustrate the problem, let $\overline{C}/K$ be a smooth projective geometrically connected curve of genus $g>0$, let $x,y\in\overline{C}(K)$ be distinct rational points, and let $C:=\overline{C}\backslash\{x,y\}$. Then $\operatorname{Pic}_{\overline{C}/K}$ is an extension of $\mathbf{Z}$ by an abelian variety, the Jacobian $J$ of $\overline{C}$. Because $\operatorname{Pic}(\overline{C}\times T)\twoheadrightarrow\operatorname{Pic}(C\times T)$ for every smooth (or even regular) $K$-scheme $T$, one can conclude that $\operatorname{Pic}^{+}_{C/K}$ is the quotient of $J$ by the point $j$ represented by the divisor $[x]-[y]$. If $j$ is not a torsion point, then this quotient is not a scheme. Thus we see that an inevitable obstacle to representability comes from quotients of group schemes by subgroups generated by points of infinite order.
We may now state the main result of the present paper. For a commutative, locally finite type $K$-group scheme $G$, we denote by $G^{t}$ the preimage in $G$ of the torsion subgroup of the étale component group $G/G^{0}$.
@@
If $K$ is a field and $X$ is a regular $K$-scheme of finite type, then there is an isomorphism $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$, where $G$ is a smooth commutative $K$-group scheme with finitely generated component group and $E\subset G^{t}$ is an étale $K$-group scheme such that $E(K_{s})$ is a free abelian group of finite rank. Furthermore, $G^{0}$ is unique up to unique isomorphism.
For results on the structure of the group $G^{0}$ appearing in Theorem 1.3, see §6. As we shall see, the most technically demanding case of Theorem 1.3 is when ${\rm{char}}(K)=p>0$. Indeed, when ${\rm{char}}(K)=0$ one may entirely skip §§34. Let us remark, however, that even for algebraically closed $K$ of characteristic $p$, these sections are required, because even though $K$ is perfect, function fields of varieties over $K$ are typically not.
Theorem 1.3 is nice, but what we would really like are results asserting representability by a scheme. Using the above theorem, we will deduce the following.
@@
Let $K$ be a field and $X$ a regular $K$-scheme of finite type with dense smooth locus. Assume that any map from $X$ into an abelian variety $A$ factors through a finite subscheme of $A$. Then $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth commutative $K$-group scheme with finitely generated component group. Furthermore, $(\operatorname{Pic}^{+}_{X/K})^{0}$ is unipotent.
@@
Abelian varieties do not admit nonconstant maps from $\mathbf{P}^{1}$ by [Mil, Prop. 6.1], hence also not from rationally connected varieties. So $X$ admits no nonconstant map to an abelian variety over $\overline{K}$, hence also not over $K_{s}$. Now apply Theorem 1.4. ∎
@@
Among the examples falling under the rubric of Theorem 1.6 are smooth connected unipotent groups. These may indeed have positive-dimensional restricted Picard functors when $K$ is imperfect. For general and specific classes of examples, see, for instance, [KMT, Th. 6.7.9] and [Ros2, Props. 5.2, 5.4, 5.5, Ex. 5.8].
@@
The problem, of course, is that we do not know whether $X$ admits a regular compactification. Even when $K$ is algebraically closed, in positive characteristic the existence of a regular compactification remains a formidable open problem. Fortunately, however, there is a weaker form of resolution which is adequate for many purposes: de Jong’s alterations theorem [deJ1, Th. 4.1]. (Actually, we will require a somewhat more precise result from [deJ2].) The difference between an alteration and a resolution is that we relax the requirement that the resolution be a birational map and only require that it be generically finite. We may replace $X$ by a dense open subscheme and thereby assume that the alteration $X^{\prime}\rightarrow X$ is finite flat. We then use faithfully flat descent to describe a line bundle on $X$ in terms of a descent datum for a line bundle on $X^{\prime}$. This is helpful because $X^{\prime}$ admits a regular compactification and therefore has “almost-representable” restricted Picard functor in the sense of Theorem 1.3.
The key difficulty in showing that the functor whose points consist of a line bundle plus descent datum on $X^{\prime}$ is (almost) representable is to prove that the descent data which arise in this manner are “uniformly bounded” in a suitable sense. To illustrate the difficulty, consider the situation in which $X=\mathbf{A}^{1}_{K}$ and $X^{\prime}=\mathbf{A}^{1}_{K^{\prime}}$ for some finite extension field $K^{\prime}/K$. (Of course, this is in a sense a silly example because $\operatorname{Pic}^{+}_{X/K}=0$ in this case, as one may see directly. But it illustrates the point most simply – without some of the technical window dressing – and so the author hopes that the reader will indulge him and pretend that he does not realize this already.) A line bundle on $X^{\prime}$ is then automatically trivial, and a descent datum for $X^{\prime}_{T}/X_{T}$ consists of nothing other than a unit on $(X^{\prime}\times_{X}X^{\prime})_{T}=(\mathbf{A}^{1}_{T})_{K^{\prime}\otimes_{K}K^{\prime}}$ (which is to say, an automorphism of the trivial line bundle) that satisfies the cocycle condition. The problem is that $K^{\prime}\otimes_{K}K^{\prime}$ may well be non-reduced when $K^{\prime}/K$ is not separable. Therefore a unit on $(\mathbf{A}^{1}_{K^{\prime}\otimes_{K}K^{\prime}})_{T}$ may have arbitrarily high degree as a “polynomial” in $\Gamma(T_{K^{\prime}\otimes_{K}K^{\prime}},\mathcal{O})[Y]$ on $\mathbf{A}^{1}$. (We put the word polynomial in scare quotes because when $T$ is not quasi-compact, a global section of $(\mathbf{A}^{1}_{K^{\prime}\otimes_{K}K^{\prime}})_{T}$ may have infinite degree.) Without bounding this degree a priori, we would obtain not representability by a scheme, but by an ind-scheme, because the set of all polynomials is not represented by a scheme. Thus we require some sort of bound on the degree. This will follow from the cocycle condition, which says that our unit $u$ satisfies
@@
The structure of this paper is as follows. In §2 we descend the almost-representability of the restricted Picard functor through finite Galois covers. In §3 we discuss the notion of boundedness for a collection of descent data, and show how boundedness of the collection of all possible descent data on certain line bundles implies the almost-representability of the restricted Picard functor. In §4 we use the notion of boundedness to construct the tools required to descend almost-representability of the restricted Picard functor through finite radicial covers. In §5 we prove the almost-representability of the restricted Picard functor for regular $K$-schemes of finite type. Finally, in §6, we study the structure of the restricted Picard functor.
1.1 Acknowledgements
I thank Kęstutis Česnavičius and the anonymous referees for many helpful comments, suggestions, and errata which have greatly improved the paper.
@@ 2 Descent through Galois covers
The uniqueness of ${\rm{Units}}_{X/K}$ follows from Yoneda’s Lemma, as does the fact that it is a commutative $K$-group scheme. To prove existence and the other assertions, we are free by Galois descent to replace $K$ by a finite Galois extension. First assume that $X$ is separated, and choose a normal compactification $\overline{X}$ of $X$. (First use the Nagata compactification theorem [Stacks, Tag 0F41] to obtain a compactification, and then normalize.) Let $A:=\Gamma(\overline{X},\mathcal{O}_{\overline{X}})$, a finite $K$-algebra. We have an inclusion $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})\hookrightarrow f_{*}{\mathbf{G}}_{\rm{m}}$. We are free to replace $K$ by a finite Galois extension, so we may assume that all of the codimension-$1$ boundary components $D_{1},\dots,D_{n}$ of $\overline{X}\backslash X$ are geometrically irreducible. For any smooth $K$-scheme $T$, $X_{T}$ is still normal, and a uniformizer of $D_{i}$ is also a uniformizer of $(D_{i})_{T}$. We thus obtain a homomorphism
@@
given by taking valuations along each of the $D_{i}$, where $f_{\operatorname{\acute{e}t}}$ is the map on étale sites induced by $f$. Then $\ker(\phi)(T)=(\overline{f}_{*}{\mathbf{G}}_{\rm{m}})(T)$ for every smooth $T/K$, where $\overline{f}\colon\overline{X}\rightarrow{\rm{Spec}}(K)$ is the structural morphism. If $A:=\Gamma(\overline{X},\mathcal{O}_{\overline{X}})$, a finite $K$-algebra, then $\overline{f}_{*}{\mathbf{G}}_{\rm{m}}=\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$. Replacing $K$ by a finite Galois extension, we may assume that $\operatorname{im}(\phi)$ is constant and that every $K$-point in $\operatorname{im}(\phi)$ lifts to a point of $(f_{\operatorname{\acute{e}t}})_{*}{\mathbf{G}}_{\rm{m}}$. So there exist units $u_{1},\dots,u_{n}$ on $X$ whose images freely generate $\operatorname{im}(\phi)$. Then every unit on $X_{T}$ for smooth connected $T$ is in the free subgroup generated by the $u_{i}$ times a unit extending over $\overline{X}_{T}$, which is to say, an element of $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})(T)$. Thus we obtain a smooth $K$-group scheme ${\rm{Units}}_{X/K}$ which is an extension of a free finite rank constant group scheme by $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$, which computes the units of $X_{T}$ for smooth $T$, and which is a subfunctor of the functor of units (because $u_{i}$ and points of $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$ define units on $X_{T}$ for any $K$-scheme $T$). This completes the proof when $X$ is separated.
Now consider the general, possibly non-separated case. Let $U\subset X$ be a separated dense open subscheme (for instance, a dense affine open). Replacing $K$ by a finite Galois extension if necessary, we may assume that all of the boundary components of $X\backslash U$ are geometrically irreducible. Let $D_{1},\dots,D_{m}$ be these components. For any smooth scheme $T$, $X_{T}$ is still normal, and a uniformizer of $D_{i}$ is also a uniformizer of $(D_{i})_{T}$. We thus obtain a homomorphism of group schemes
@@
given by taking valuations along each of the $D_{i}$, and $\ker(\phi)(T)=(f_{*}{\mathbf{G}}_{\rm{m}})(T)$ for every smooth $T/K$. Because the codomain of $\phi$ is étale, $\ker(\phi)^{0}={\rm{Units}}_{U/K}^{0}$ and $\ker(\phi)$ is smooth. Furthermore, the algebraic subrings of $\Gamma(X,\mathcal{O}_{X})$ and $\Gamma(U,\mathcal{O}_{U})$ agree by normality of $X$. Thus $\ker(\phi)^{0}=\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$, hence $\ker(\phi)^{0}$ defines a subfunctor of $f_{*}{\mathbf{G}}_{\rm{m}}$. A similar argument to the one earlier involving the group freely generated by the $u_{i}$ then shows that $\ker(\phi)$ is a subfunctor of $f_{*}{\mathbf{G}}_{\rm{m}}$, hence ${\rm{Units}}_{X/K}=\ker(\phi)$. That the component group is free of finite rank follows from the analogous assertion for $U$. ∎
@@
For a commutative $K$-group scheme $M$ of finite type equipped with an action by the finite group $G$, $\mathscr{H}^{i}(G,M)$ is represented by a smooth commutative $K$-group scheme of finite type, which is affine if $M$ is so.
@@
First, nothing changes if we replace $M$ by its maximal smooth $K$-subgroup scheme, so we may assume that $M$ is smooth. The functor $\mathcal{Z}^{n}(G,M)$ of $n$-cocycles $G^{n}\rightarrow M$ is readily checked to be representable by the closed subscheme of $M^{G^{n}}$ corresponding to the cocycle condition. We also have coboundary maps $d_{n+1}\colon M^{G^{n}}\rightarrow\mathcal{Z}^{n+1}(G,M)$. Because $M^{G^{n}}$ is smooth, the quotient group sheaf $Q^{i}:=\mathcal{Z}^{i}(G,M)/{\rm{im}}(d_{i+1})$ on the smooth-étale site of $K$ agrees with the usual fppf quotient group scheme. (That is, the étale quotient is already a scheme by [Stacks, Tag 0B8G], hence an fppf sheaf.) Replacing $Q^{i}$ by its maximal smooth $K$-subgroup scheme ([CGP, Lem.C.4.1, Rem. C.4.2]) does not change the underlying functor on the smooth-étale site of $K$, so we see that $\mathscr{H}^{i}(G,M)$ is represented by a smooth commutative finite type $K$-group scheme, which is affine if $M$ is. ∎
@@
Let $M$ be a commutative $K$-group scheme with finitely generated component group, and suppose that $M$ comes equipped with an action by a finite group $G$. Then ${\mathscr{H}}^{i}(G,M)$ is (represented by) a smooth $K$-group scheme with finitely generated component group. For $i>0$, ${\mathscr{H}}^{i}(G,M)$ is finite type, and it is affine if $M^{0}$ is so.
@@
As before, we may replace $M$ by its maximal smooth $K$-subgroup scheme and thereby assume that $M$ is smooth. Let $E:=M/M^{t}$ be the torsion-free quotient of the component group of $M$. By Galois descent, we may assume that $K$ is separably closed, so that $E$ is constant and the quotient map $M\rightarrow E$ admits a section, so that $M\simeq M^{t}\times E$. We are thus reduced to the étale and finite type cases, which are handled by Propositions 2.3 and 2.4. ∎
@@
We say that a sheaf $\mathscr{F}$ of abelian groups on the smooth-étale site of ${\rm{Spec}}(K)$ is almost-representable when it is of the form $G/E$ for a smooth commutative $K$-group scheme $G$ with finitely generated component group and a subsheaf $E\subset G$ represented by an étale $K$-group scheme $E$ with $E(K_{s})$ finitely generated.
We will soon obtain a slight strengthening of the above definition that is sometimes useful (Lemma 2.9).
@@
The following lemma gives a useful strengthening of the definition of almost-representable sheaves.
@@
Let $\mathscr{F}$ be an almost-representable sheaf on the smooth-étale site of ${\rm{Spec}}(K)$. Then one has an isomorphism $\mathscr{F}\simeq G/E$, where $G$ is a smooth commutative $K$-group scheme with finitely generated component group, and $E\subset G^{t}$ is a subsheaf represented by an étale $K$-group scheme such that $E(K_{s})$ is free of finite rank.
@@
Let $\mathscr{F}\simeq H/F$ be a presentation of $\mathscr{F}$ as an almost-representable sheaf. Let $C:=H/H^{t}$ denote the torsion-free quotient of the component group of $H$, and let $\phi\colon F\rightarrow C$ denote the composition $F\subset H\rightarrow C$. First assume that $\phi(F)=0$. Then $F\subset H^{t}$. Let $T\subset F$ denote the torsion subgroup. Because $T$ is finite, $\overline{H}:=H/T$ is a smooth $K$-group scheme with finitely-generated component group. (It is a scheme by [Stacks, Tag 0B8G], and smooth by [EGA, IV1, Prop. 17.3.3].) Then $\mathscr{F}\simeq H/F=\overline{H}/(F/T)$ gives a presentation of $\mathscr{F}$ of the desired form.
Now assume that $\phi(F)\subset C$ is nonzero. Then $F$ defines an element of $\operatorname{Ext}^{1}(\phi(F),\ker(\phi))$. By Lemma 2.8, this element is killed by some $N>0$. The multiplication by $N$ map induces an isomorphism $\phi(F)\simeq[N]\phi(F)$, so the extension of group schemes